OpenStudy (anonymous):

if a=b2/3 and b=c3/4 what is the relation between a and c?

OpenStudy (anonymous):

Place equation 2 in equation 1 a = [(c3/4)*2] / 3 a = 6c / 12 a = c/2

OpenStudy (anonymous):

so a=c1/2?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

just gave it